07 8 4 0 O 10 2 O 11 12 5 13 O 14 O 15 3 16 O 17 6 018х Y 60% Question 2 Z т 10% PA 500 PB 300 4% 30% 2) A householder is considering insuring the contents of his house for 1 year
Answer #1

PLEASE LIKE THIS ANSWER, IT HELPS ME A LOT. THANK YOU

(a)

No theft 96% 7 No loss 1 theft 4% 8 Lose 10.000 TL 50% No policy 4 9 Lose 20.000 TL 30% 0 Company A 10 Lose 40.000 TL 20% No

(b)

Path to number 7: Total Cost: 0

Path to number 8: Total Cost: 10,000 TL

Path to number 9: Total Cost: 20,000 TL

Path to number 10: Total Cost: 40,000 TL

Path to number 11: Total Cost: 500 TL

Path to number 12: Total Cost: 500 TL

Path to number 13: Total Cost: 500 TL

Path to number 14: Total Cost: 500 TL

Path to number 15: Total Cost: 300 TL

Path to number 15: Total Cost: 300 + (10.000 x 0.3) = 3,300 TL

Path to number 15: Total Cost: 300 + (20.000 x 0.3) = 6,300 TL

Path to number 15: Total Cost: 300 + (40.000 x 0.3) = 12,300 TL

Consider Node 4, expected monetary value for this node will be: Total Cost = (10,000 x 0.5) + (20,000 x 0.3) + (40,000 x 0.2) = 19,000 TL

Consider Node 1, expected monetary value for this node will be: Total Cost = (19,000 x 0.04) + (0 x 0.96) = 760 TL

Consider Node 5, expected monetary value for this node will be: Total Cost = (500 x 0.5) + (500 x 0.3) + (500 x 0.2) = 500 TL

Consider Node 2, expected monetary value for this node will be: Total Cost = (500 x 0.04) + (500 x 0.96) = 500 TL

Consider Node 6, expected monetary value for this node will be: Total Cost = (3,300 x 0.5) + (6,300 x 0.3) + (12,300 x 0.2) = 6,000 TL

Consider Node 3, expected monetary value for this node will be: Total Cost = (6,000 x 0.04) + (300 x 0.96) = 528 TL

Therefore, So working with Company A would have the least expected cost. We should chose Company A

PLEASE LIKE THIS ANSWER, IT HELPS ME A LOT. THANK YOU